OpenStudy (anonymous):

Auditions for a play in New York were held with 100 men and women trying out for the lead singer position. If twice the number of men less three times the number of women is equal to 20, how many men auditioned? 44 40 76 64

OpenStudy (anonymous):

make two equation

OpenStudy (anonymous):

how do I do that?

OpenStudy (anonymous):

m + w = 100 2m - 3w = 20

OpenStudy (anonymous):

m+w =100 2m-3w=20

OpenStudy (anonymous):

alright then what?

OpenStudy (anonymous):

then multiply first equation by -2

OpenStudy (anonymous):

-50

OpenStudy (anonymous):

-2m-2w=-200 2m-3w=20

OpenStudy (anonymous):

oh I divided 100 from -2 oops

OpenStudy (anonymous):

-5w=-180 w=36 m=64

OpenStudy (anonymous):

You can use substitution: m=100-w so then the second equation can be rewritten as 2(100-w) - 3w = 20 200-2w-3w=20 200-5w=20 -5w=-180 5w=180 w=36 then you substitute that back into the first equation: m+w=100 m+36=100 m=64

OpenStudy (anonymous):

thanks guys!

OpenStudy (anonymous):

yw